OpenStudy (anonymous):

solve using matrices x+3y=13 2x-y=-9

OpenStudy (amistre64):

pull the coefficients and constants, then row reduce

OpenStudy (unklerhaukus):

\[\begin{align}x+3y&=13\\ 2x-y&=-9\end{align}\] \[\begin{bmatrix}1&3\\2&-1\end{bmatrix}\begin{bmatrix}x\\y\end{bmatrix}=\begin{bmatrix}13\\-9\end{bmatrix}\] \[\left[\begin{array}{cc|c}1&3&13\\2&-1&-9\end{array}\right]\]